OpenStudy (anonymous):

Find the derivative of f when f(x) = 2 tan(x) + 3 cot(x)

OpenStudy (cwrw238):

derivative of tan x is sec^x can't remember derivative of cot x - try googling it first term of derivative is 2 sec^2 x

OpenStudy (anonymous):

of cotx? its -csc^2

OpenStudy (cwrw238):

http://www.math.com/tables/derivatives/tableof.htm this link lists derivatives

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

sol: let y=f(x) y=2tan(x)+3cot(x) =2dy/dxtanx+3dy/dxcotx =2*sec^2x+3*-cosec^2x =2sec^2x-3cosec^2x

OpenStudy (cwrw238):

oh ok so f'(x) = 2 sec^2 x - 3 csc^2 x

OpenStudy (anonymous):

think they want me to make in ration form its number 3 on the hw i have attached

OpenStudy (anonymous):

is it answer??

OpenStudy (anonymous):

thats not one of the answer choices

OpenStudy (anonymous):

muhammand please help

OpenStudy (anonymous):

estudier please help

OpenStudy (anonymous):

are these questions???

OpenStudy (anonymous):

yes the document i have attached is the homework i am looking at number 3 specifically

OpenStudy (cwrw238):

ah what they have done is convert the sec and csc to sin and cos

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

so what is the answer

OpenStudy (cwrw238):

2 sec^2 x - 3 csc^2 x = 2 * 1/cos^2 x - 3 * 1 / sin^2 x

OpenStudy (cwrw238):